OpenStudy (anonymous):

If you want your retirement account to have a balance of $200000 when you retire in 35 years, how much should you invest now at a rate of 8% interest compounded continuously?

OpenStudy (tkhunny):

Solve for A. \[A\cdot e^{0.08t} = 200000\] Is it looking familiar?

OpenStudy (anonymous):

yes I have the equation set up as Pe^(.08*35)=200000 but im not quite sure how to solve it

OpenStudy (tkhunny):

What's to solve? Evaluate the numbers. Hvwe youan exponent function on your calculator? You'll need one. I suppose you could rewrite a little, first: \[P = 200000\cdot e^{-0.08\cdot 35}\]

OpenStudy (anonymous):

ooooooo ok thank you,,, so you just multiply the exponential function over, wouldnt you need to divide it though to get in on the other side?

OpenStudy (tkhunny):

I did divide, but the fraction was cumbersome so I rewrite it with the negative exponent. You need to get natural and fluent at this.

OpenStudy (anonymous):

ooooo bc the exponent is negative, makes it division?

OpenStudy (tkhunny):

\[\frac{1}{B} = B^{-1}\] Were you paying attention in algebra class? :-)

OpenStudy (anonymous):

i took algebra class like 10 years ago but thx

OpenStudy (tkhunny):

Fair enough. That is a valid excuse. You will need it, though. Get it up to speed and good work!